# Infra-red thermographic inversion in ST40

### v1

@@ Discussion
The goal of this paragraph is to explain the reasons behind the behaviour of $\varepsilon_{*}(\mathcal{E})$ of Sec. 4.1.2, with $\Delta x=\Delta y$. With the inequality $o(\Delta t)\ll o(\Delta x^{2})$ ensured in the space convergence (Sec. 3.3.2) and the convergence properties of Eq. (7) and Eq. (8), the error $\varepsilon_{*}(\mathcal{E})$ from Eq. (12) must abide to a first order convergence in space. This holds by virtue of the ordering $o(\Delta t)\ll o(\Delta x^{2})\sim o(\Delta x^{2})+o(\Delta y^{2})\ll o(\Delta y)\ll 1$.
On the contrary, an explicit dependence on $\Delta t$ is almost totally absent in Table 1 (last column). Granted $o(\Delta x^{2})\ll o(\Delta t)$ in the time convergence (Sec. 3.3.2), this suggests $o(\Delta x^{2})+o(\Delta y^{2})\sim o(\Delta x^{2})\ll o(\Delta t)\lesssim o(\Delta y)\ll 1$ to be in place, where $o(\Delta t)\lesssim o(\Delta y)$ implies that the trend in time of $\varepsilon_{*}(\mathcal{E})$ is hampered by $o(\Delta y)$.
In other words, the first-order error $o(\Delta y)$ affecting the heat flux $q_{i}^{n}$ via Eq. (5) is always the dominant contribution, with regards to energy conservation. Although the resulting error is still acceptable (Sec. 4.1.2), a point of improvement for FAHF would be using higher-order discretisations in the fashion of Ref. [49]’s instead of Eq. (5), were higher precision required.
@@
As also suggested by Ref. [14], Sec. 4.1.4 clarifies how a precise heat flux can be calculated by the algorithm with a proper choice of $\Delta y<\Delta x$, without the need of an extremely small $\Delta x$. This hence allows for saving computational cost.
However, the assessment of the solution $q_{i}^{n}$ and its Eich fit in terms of $\Delta x$ (potentially with $\Delta x<\Delta x_{\text{IR}}$), as performed here, is also recommended. For small enough $\Delta y$, $\Delta x$ becomes the bottleneck in the error $o(\Delta x^{2})+o(\Delta y)$ of Eq. (5), even more so if higher-order discretisations with $\Delta y$ are employed.
@@
The fact that the resolution along $y$ (independent on the diagnostic apparatus) mostly determines the precision of $q_{i}^{n}$ (within the limits of Sec. 5.2) does loosen the constraints on the required IR camera resolution $\Delta x_{\text{IR}}$ (Sec. 2.2).
The suitable $\Delta x_{\text{IR}}$ is such that to allow capturing features of $T_{\text{IR}}(x)$, a smoother and broader profile than its corresponding $q_{i}^{n}$. Instead, the suitable $\Delta x$ follows from Sec. 5.2. The two resulting figures can therefore differ, with $\Delta x_{\text{IR}}\geq\Delta x$, and potentially $\Delta x_{\text{IR}}\gg\Delta x$.
Nevertheless, the recurring diagnosis of power fall-off lengths significantly in defect of the Eich scaling predictions [57] does rise the concern on the resolution $\Delta x_{\text{IR}}$ required by IR systems in future devices.
Notably, the deviation of the FAHF result from the ideal Eich fit in $x\in[10;20]$ mm $\forall\Delta y\leq 1.3\cdot 10^{-4}$ m in Fig. 5a is a genuine feature, because $\Delta x=5.5\cdot 10^{-4}\;\text{m}$ adopted in Sec. 4.1.4 is above the IR camera resolution $\Delta x_{\text{IR}}=2.1\cdot 10^{-4}$ m (Sec. 2.2). The effective spatial resolution of the ST40 IR system does allow for the appreciation of fine-scale structures in the heat flux footprint, not dissimilar for those predicted by the XGC code[48, 58].
@@
Although Sec. 4.2.1 already concluded on the appropriateness of Asm. #3 in FAHF, further feature worth investigating appear from the comparison.
@@
The error FAHF vs. COMSOL3D of $9.6\%\pm 1.4\%$ in $[40;90]$ ms is greater than the $6.9\%\pm 1.1\%$ in $[90;140]$ ms by a statistically significant margin. The same is not true for FAHF vs. COMSOL2D instead, for which $4.6\%\pm 1.0\%$ in $[40;90]$ ms and $4.8\%\pm 1.6\%$ in $[90;140]$ ms. This 3D-related feature likely follows from the appearance of the strike point in the camera FoV during $[90;140]$ ms, which leads to $|\partial_{x}T_{\text{IR}}|\gg|\partial_{\phi}T_{\text{IR}}|$ (Sec. 3.1.2), i.e. a stronger tendency for the heat flux to flow poloidally (favouring a 2D treatment) rather than toroidally (inviting 3D flows).
Additionally, COMSOL3D maximum is always in defect of its COMSOL2D counterpart all throughout $[40;140]\;\text{ms}$. COMSOL3D hence predicts that the same PFC temperature (i.e. $T_{\text{IR}}$) tends to be reached by means of a lower surface heat flux. This effect is again tied to the three-dimensionality of the problem, and presumably due to the additional heat flux that flows along $\phi$ (i.e. across the chords, Sec. 3.1.2) from $k<10$ towards the reference chord $k=10$ in COMSOL3D (and in reality), by virtue of $\partial_{\phi}T_{\text{IR}}<0$ (Fig. 1b). This additional contribution along $\phi$, accounted neither by FAHF nor by COMSOL2D, helps rising the temperature along chord 10 in the 3D case, thus requiring a lower $q_{i}^{n}$ along $y$.
Moreover, because of the FAHF incapability of resolving rapid dynamics, FAHF errors diverge towards 100% during the fast transients at the instants $t\sim 10$ ms and $t\sim 160$ ms, over a time of $\sim 5$ ms. This is a thermal information which, in the legitimate approximation of semi-infinite solid [12] (Sec. 3.1.1), affects a region of thickness $2\sqrt{\alpha\cdot 5\;\text{ms}}\sim 1$ mm along $y$. According to the resistance-capacitor (RC) circuital interpretation[12], this information then exponentially dissipates with a time constant in the order of $(1\;\text{mm})^{2}/\alpha\sim 19$ ms. And the FAHF error is indeed seen to lose memory of the information over a timescale of this magnitude, following the $t\sim 10$ ms and $t\sim 160$ ms events. The exponential fit of the neat FAHF vs. COMSOL3D error over the time window $[10;35]$ ms returns a time constant of 12 ms with a satisfactory $R^{2}=98\%$. Notably, this implies that the error during the initial phase ($t\sim 10$ ms) does not detrimentally impact the longer-term dynamics predicted by FAHF.
Conversely, the error at $t\gtrsim 160$ ms, despite firstly dropping, remains significant. This is presumably implied by the nature of this phase, an inertial cooling after the plasma magnetic topology reverts back to limited (with a sudden drop of divertor heating). Such a simple dynamics of inertial cooling, especially in conjunction with adiabatic boundaries (Sec. 3.1.4), is highly sensitive to the conditions at its beginning ($t\sim 160$ ms), which are very different between FAHF and COMSOL ($\varepsilon^{n}(q)\sim 100$%).
@@
The FAHF vs. COMSOL3D error in space $\varepsilon_{i}(q)$ increases (i) towards the toroidal-running gaps, i.e. $x=0$ and $x=32$ mm. This is expected because of the first-order boundary conditions of FAHF (Sec. 3.1.4), and because of the higher degree of three-dimensionality necessarily involved around these regions: the adiabatic boundary invites the heat to flow elsewhere than along $x$, e.g. along $\phi$, not modelled in 2D. However, $\varepsilon_{i}(q)$ is also enhanced (ii) around the centre of the tile, i.e. $x\sim 15$ mm. This can be ascribed to the dual of (i): the centre of the tile allows for a 3D heat flux behaviour by offering the loosest constraints, by virtue of being the point farthest away from any adiabatic boundary. A gradual transition towards 2D between these 3D extremes (i) and (ii) instead characterises the remaining regions, where the error $\varepsilon_{i}(q)$ drops.
@@
Negative surface heat fluxes are here predicted in all the three cases: (i) in $(0;10]$ ms in Fig. 6c; (ii) at $t_{*}=128$ ms on the left of the peak in Fig. 6d, $x\in[0;5]$ mm; (iii) all throughout the the divertor cooling phase (Sec. 5.4.1) at $t\gtrsim 160$ ms, not shown.
With regards to (i), vibrations during the plasma start-up are the most reasonable explanation. The same does not instead hold for (ii), presumably affected by the presence of an impurity surface layer and non-negligible black-body emission. Last, the negative heat flux in (iii), comparable in magnitude to the maximum during such phase ($0.5\;\text{MW}\cdot\text{m}^{-2}$), is entirely imputable to the black-body emission during the inertial cooling. Both the above-mentioned points are worthwhile further investigating in the future, despite not impacting the conclusions of the present study.
@@
Data of divertor surface temperature are recorded by ST40’s IR camera system with the highest effective spatial resolution available worldwide ($0.21\pm 0.03$ mm/pixel). Leveraging such highly-resolved data, FAHF, the new Tokamak Energy’s tool for inverse thermographic analyses, is developed as a finite-difference explicit solver of the unsteady heat conduction equation. Under the simplifying assumptions of temperature-independent material properties, single molybdenum layer instead of multi-material multi-layered PFCs, and two-dimensionality, FAHF calculates the surface heat flux impinging on the divertor.
Answering the question of whether FAHF qualifies as a suitable tool for inverse thermographic analyses is the focus of this work. And the answer is a positive one.
The formal verification of FAHF by means of time and space convergence is successful, in fact showing the expected first- and second-order convergence with $\Delta t$ and $\Delta x=\Delta y$ (for the sake of simplicity), respectively. The most appropriate combination of $\Delta t$ and $\Delta x$ is chosen accordingly, such as to guarantee a satisfactory compromise between precision of the temperature output and computational cost.
@@
The behaviour of the surface heat flux (along the divertor depth $y$), leading metric in any inverse thermographic analyses, is also thoroughly scrutinised. The Eich fit [6] is used to extract physical parameters of interest. The two separate cases of $\Delta x=\Delta y$ and $\Delta x\neq\Delta y$ are investigated. On the one hand, in both instances a strong sensitivity of the fitting parameters as a function of the spatial resolution is found, with variations up to a factor 2.5. On the other hand, the two instances ultimately show a convergence of the fitting parameters towards common values, as the spatial resolution is enhanced. Converged values of the $\Delta x\neq\Delta y$ case are achieved at a lower computational cost, by virtue of the leading dependence of the surface heat flux being $\Delta y$ (divertor depth), rather than $\Delta x$ (along the divertor). Although this dependence is commonly leveraged in inverse thermographic analyses [14, 49], the error implied by $\Delta x$ may become noticeable in the limit $\Delta y\ll\Delta x$, and detrimentally affect the surface heat flux. Routinely analysing the surface heat flux as a function of the space step $\Delta x$ along the divertor is thus recommended.
With the above guaranteeing a satisfactory precision in FAHF, the comparison with the renowned COMSOL Multiphysics® is well-posed, and evinces a satisfactory accuracy of FAHF in any instance analysed.
@@
The average agreement within $8.3\%$ in time and $10\%$ in space among FAHF and COMSOL (3D) demonstrates the legitimacy of the 2D treatment adopted in FAHF. Furthermore, the simplistic implementation in FAHF of temperature-independent material properties and single molybdenum layer is proven to be appropriate in ST40 shots: the corrections implied by including these features in COMSOL (2D) are shown to be of minor entity ($<1\%$).
@@
Regions of disagreement in time among FAHF and COMSOL are circumscribed to the beginning/end of the plasma shot, and at the termination of the divertor heating phase as the plasma abruptly reverts to being limited. These time windows of fast dynamics could be properly resolved by improving the temporal discretisation properties of FAHF [14], but are outside the scope of the present work, and outside the time window of interest for thermographic studies.
@@
Disagreement in space among FAHF and COMSOL ($\sim 20\%$) is instead relegated to the regions in proximity of the toroidal-running gaps between divertor tiles. This is true in the 2D case, presumably because of the low-precision Neumann boundary conditions implemented in FAHF, and in the 3D case, also because of the adiabatic nature of these boundaries which invites a toroidal flow of the heat flux. Either instance does not pose a threat to the heat flux calculated by FAHF: the above-mentioned areas are also affected by further physical effects (e.g. tile side-heating due to finite ion Larmor radius [46]) which are not included in any thermographic inversion, and must therefore be under-emphasised anyway.
@@
In conclusion, Tokamak Energy’s FAHF does qualify as a trustworthy tool, internally consistent and capable of providing a satisfactory precision and accuracy at an affordable computational cost, modulo a proper choice of $\Delta t$, $\Delta x$ and $\Delta y$ which this work identifies: $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m.
@@
In perspective, further studies are planned in ST40 with the two-fold goal of improving the IR thermographic inversion experimentally and numerically. The current IR camera acquisition range has already been extended from $[20;118]$ to $[20;200]$, and a detailed cross-comparison with Langmuir probe measurements is underway. The installation of a second IR camera is being considered, for the exploitation of a dual-band adaptor [59] would allow ameliorating any dependence on the uncertain surface emissivity. The ongoing commissioning of the dedicated heat-flux testing facility in Tokamak Energy is ongoing, and will pave the way for a validation of FAHF in a controlled environment. Improvements to the numerics of FAHF to handle fast dynamics and impurity surface layers [14] are also taken into account.

### v2

@@ Discussion
The goal of this paragraph is to discuss the behaviour of $\varepsilon_{*}(\mathcal{E})$ of Sec. 5.1.2, with $\Delta x=\Delta y$. The underlying algebra is explicitly detailed in Appendix Algebra of the energy balance error, and states how the error $\varepsilon_{*}(\mathcal{E})$ must abide to a first order convergence in space, the cases of Table 3 being considered.
On the contrary, an explicit dependence on $\Delta t$ is almost totally absent in Table 3 (last column) suggesting that $o(\Delta t)\lesssim o(\Delta y)$, i.e. the trend in time of $\varepsilon_{*}(\mathcal{E})$ is hampered by $o(\Delta y)$.
The first-order error $o(\Delta y)$ affecting the heat flux $q_{i}^{n}$ via Eq. (9) is indeed found to be the dominant contribution, with regards to energy conservation. Although the resulting energy error is still acceptable (Sec. 5.1.2), a point of improvement for FAHF would be using higher-order discretisations in the fashion of Refs.[58, 14]’s instead of Eq. (9), were higher precision in energy required.
@@
As also suggested by Ref.[18], Sec. 5.1.4 clarifies how a precise heat flux can be calculated at a lower cost with a proper choice of $\Delta y<\Delta x$, without the need of an extremely small $\Delta x$.
However, the assessment of the error in $q_{i}^{n}$ and in the energy balance (Sec. 5.1.2) as a function of $\Delta x$ is also recommended. For small enough $\Delta y$, $\Delta x$ becomes the bottleneck in $o(\Delta x^{2})+o(\Delta y)$ of Eqs. (9) and (15), potentially even more so if higher-order discretisations of $q_{i}^{n}$ with $\Delta y$ are employed.
From the considerations above, the ultimate $\Delta x$ selected might fall below the spatial resolution of the camera, i.e. $\Delta x<\Delta x_{\text{IR}}$. This is not of concern as long as a linear interpolation of $T_{\text{IR}}$ is employed (Sec. 4.1.4). If, on the one hand, $\Delta x<\Delta x_{\text{IR}}$ does not add any physics-relevant information, on the other hand, it plays a role in the numerics of the problem. And the cautious user should not be restrained by $\Delta x_{\text{IR}}$ when identifying the most proper $\Delta x$.
@@
The fact that the resolution along $y$ (independent on the diagnostic apparatus) mostly determines the precision of $q_{i}^{n}$ does loosen the constraints on the required IR camera resolution $\Delta x_{\text{IR}}$ (Sec. 2.2.1).
While $\Delta x$ follows from Sec. 6.2, the suitable $\Delta x_{\text{IR}}$ is such that to allow capturing features of $T_{\text{IR}}$, a smoother and broader profile than the corresponding surface heat flux.
Nevertheless, power fall-off lengths significantly narrower than the Eich scaling predictions [37, 38, 39] do rise the concern on the resolution $\Delta x_{\text{IR}}$ required by IR thermographic systems in current and future devices. This clearly reverberates on the choice of $\Delta x$ and $\Delta y$ (Sec. 6.2).
Last, the FAHF result $q_{i}^{n}$ is seen to very closely follow $q_{\text{Eich}}\;\forall\Delta x\geq 5.5\times 10^{-4}$ m in Fig. 5a. Instead, a deviation from the Eich fit in $x\in[10;20]$ mm is noticeable $\forall\Delta x\leq 2.7\times 10^{-4}$ m, and also predicted by COMSOL in Fig. 6b. This deviation is a genuine feature because $\Delta x>\Delta x_{\text{IR}}=2.1\times 10^{-4}$ m (Sec. 2.2.1). The effective spatial resolution of the ST40 IR system does allow for the appreciation of fine-scale structures in the heat flux footprint, not dissimilar for those predicted by the XGC code[54, 68], meritorious of further studies.
@@
Although Sec. 5.2.1 already concluded on the appropriateness of Asm. #3 in FAHF, further investigations are warranted in the light of Sec. 4.2.2. Only CH#2 of Sec. 4.3.1, most suitable for a 2D analysis (Sec. 5.2.1).
@@
The error FAHF vs. COMSOL3D of $10\%$ in $[40;90]$ ms is greater than the $7\%$ in $[90;140]$ ms. The same is not true for the error FAHF vs. COMSOL2D instead, which reads $5\%$ in both $[40;90]$ ms and $[90;140]$ ms. This 3D-related feature likely follows from the appearance of the strike point in view of the camera during $[90;140]$ ms, which leads to $|\partial_{x}T_{\text{IR}}|\gg|\partial_{\phi}T_{\text{IR}}|$ (Sec. 4.1.2), i.e. a stronger tendency for the heat flux to flow poloidally (favouring a 2D treatment, hence lowering the error) rather than toroidally (inviting 3D flows, thus enhancing the error).
Additionally, COMSOL3D maximum is systematically in defect of its COMSOL2D counterpart all throughout $[40;140]\;\text{ms}$. COMSOL3D hence predicts that the same PFC temperature (i.e. $T_{\text{IR}}$) tends to be reached by means of a lower surface heat flux along $y$. This effect is again tied to the three-dimensionality of the problem, and presumably due to the additional heat flux that flows along $\phi$ (i.e. across the chords, Sec. 4.1.2) towards the CH#2 in COMSOL3D (and in reality), by virtue of $\partial_{\phi}T_{\text{IR}}<0$ (Fig. 1b). This additional contribution along $\phi$, accounted neither by FAHF nor by COMSOL2D, helps rising the temperature along CH#2 in the 3D case, thus requiring a lower $q_{i}^{n}$ along $y$.
Moreover, because of the FAHF incapability of resolving rapid dynamics, FAHF errors diverge towards 100% during the fast transients at the instants $t\sim 10$ ms and $t\sim 160$ ms, over a time of $\sim 5$ ms. This error is a thermal information which, in the legitimate approximation of semi-infinite solid [10] (Sec. 4.1.1), affects a region of thickness $2\sqrt{\alpha\times 5\;\text{ms}}\sim 1$ mm along $y$. According to the resistance-capacitor circuital interpretation[10], this information then exponentially dissipates with a time constant in the order of $(1\;\text{mm})^{2}/\alpha\sim 19$ ms. And the FAHF error is indeed seen to drop over a timescale of this magnitude, following the $t\sim 10$ ms and $t\sim 160$ ms events. The exponential fit of the neat FAHF vs. COMSOL3D error over the time window $[10;35]$ ms returns a time constant of 12 ms with a satisfactory $R^{2}=98\%$. This crucially implies that the error during the initial phase ($t\sim 10$ ms) does not detrimentally impact the longer-term dynamics predicted by FAHF (loss of memory).
On the contrary, the error at $t\gtrsim 160$ ms, despite firstly dropping, remains significant. This is presumably implied by the nature of this phase, an inertial cooling after the plasma magnetic topology reverts back to limited (with a sudden drop of divertor heating). Such a simple dynamics of inertial cooling, especially in conjunction with adiabatic boundaries (Sec. 3.1), is highly sensitive to the conditions at its beginning ($t\sim 160$ ms), which are very different between FAHF and COMSOL ($\varepsilon^{n}(q)\sim 100$%).
@@
The FAHF vs. COMSOL3D error in space $\varepsilon_{i}(q)$ increases (i) towards the toroidal-running gaps, i.e. $x=0$ and $x=32$ mm. This is expected because of the first-order boundary conditions of FAHF (Sec. 4.1.4), and because of the higher degree of three-dimensionality necessarily involved around these regions: the adiabatic boundary invites the heat to flow elsewhere than along $x$, e.g. along $\phi$, not modelled in 2D. Because further physical effects thereabouts occurring (e.g. tile side-heating due to finite ion Larmor radius [48, 69]), however not included in any thermographic inversion, (i) must be under-emphasised anyway, and is therefore not of concern.
Nonetheless, $\varepsilon_{i}(q)$ is also enhanced (ii) around the centre of the tile, i.e. $x\sim 15$ mm. This can be ascribed to the dual of (i): the centre of the tile allows for a 3D heat flux behaviour by offering the loosest constraints, by virtue of being the point farthest away from any adiabatic boundary. A gradual transition towards 2D between these 3D extremes (i) and (ii) instead characterises the remaining regions, where the error $\varepsilon_{i}(q)$ drops.
@@
Negative surface heat fluxes are here predicted in all the three cases: (i) in $(0;10]$ ms in Fig. 6a; (ii) at $t_{*}=128$ ms on the left of the peak in Fig. 6b, $x\in[0;5]$ mm; (iii) all throughout the the divertor cooling phase (Sec. 6.4.1) at $t\gtrsim 160$ ms, not shown.
With regards to (i), vibrations during the plasma start-up are the most reasonable explanation. The same does not instead hold for (ii) nor (iii), presumably affected by the presence of an impurity surface layer and non-negligible black-body emission. Both the above-mentioned points are worthwhile further investigating in the future, despite not impacting the conclusions of the present study.
@@
Data of divertor surface temperature are recorded by ST40’s IR camera system with the highest effective spatial resolution available worldwide ($0.21\pm 0.03$ mm/pixel). Leveraging such highly-resolved data, FAHF, the new Tokamak Energy’s tool for inverse thermographic analyses, is developed as a finite-difference explicit solver of the unsteady heat conduction equation. Under the simplifying assumptions of temperature-independent material properties, all-molybdenum mono-layered PFCs, and two-dimensionality, FAHF calculates the surface heat flux impinging on the divertor.
In this work, we have demonstrated the suitability of FAHF as a tool for inverse thermographic analyses.
The formal verification of the code by means of time and space convergence is successful, in fact showing the expected first- and second-order convergence with $\Delta t$ and $\Delta x=\Delta y$ (for the sake of simplicity), respectively. The most appropriate combination of $\Delta t$ and $\Delta x$ is chosen accordingly, such as to guarantee a satisfactory compromise between precision of the temperature output and computational cost.
@@
The behaviour of the surface heat flux (along the divertor depth $y$) is also thoroughly scrutinised. The Eich fit is used to extract physical parameters of interest in the two separate cases of $\Delta x=\Delta y$ and $\Delta x\neq\Delta y$. On the one hand, in both instances a strong sensitivity of the fitting parameters as a function of the spatial resolution is found, with variations up to a factor 2.5. On the other hand, the two instances ultimately show a convergence of the fitting parameters towards common values, as the spatial resolution is enhanced. Converged values of the $\Delta x\neq\Delta y$ case are achieved at a lower computational cost, by virtue of the leading dependence of the surface heat flux being $\Delta y$ (divertor depth), rather than $\Delta x$ (along the divertor). Although this dependence is commonly leveraged in inverse thermographic analyses, the error implied by $\Delta x$ may become noticeable in the limit $\Delta y\ll\Delta x$, and detrimentally affect the surface heat flux. Routinely analysing the surface heat flux as a function of the space step $\Delta x$ along the divertor is thus recommended. This has consequences in dictating the appropriate spatial resolution of the IR thermographic system, which this study comments.
@@
With the above guaranteeing a satisfactory precision of FAHF, the comparison with the renowned COMSOL Multiphysics® is well-posed, and evinces a satisfactory accuracy of FAHF. The average agreement within $8\%$ in time and $10\%$ in space among FAHF and COMSOL (3D) demonstrates the legitimacy of the 2D treatment adopted in FAHF. However, the agreement in space particularly worsens in the proximity of the poloidal-running gaps, and the resulting 22% should discourage the application of a 2D treatment in such areas. The simplistic implementation in FAHF of temperature-independent material properties and single molybdenum layer is proven to be appropriate in ST40 shots: the corrections implied by including these features in COMSOL (2D) are shown to be of minor entity ($<1\%$).
@@
Regions of inaccuracy in time of FAHF are circumscribed to the beginning/end of the plasma shot, and at the termination of the divertor heating phase. These fast dynamics could be properly resolved by improving the temporal discretisation properties of FAHF, but are outside the scope of the present work, outside the time window of interest for thermographic studies, and outside the time resolution of the system.
@@
In space, the error $\sim 20\%$ of FAHF is instead relegated to the regions in proximity of the toroidal-running gaps between divertor tiles. This is true in the 2D case, presumably because of the low-precision Neumann boundary conditions implemented in FAHF, and in the 3D case, also because of the adiabatic nature of these boundaries which invites a toroidal flow of the heat flux. Either instance does not pose a threat to the heat flux calculated by FAHF, as the above-mentioned areas are to be under-emphasised anyway.
@@
In conclusion, Tokamak Energy’s FAHF does qualify as a trustworthy tool, internally consistent and capable of providing a satisfactory precision and accuracy at an affordable computational cost, modulo a proper choice of the inputs which this work identifies: $\Delta t=7.2\times 10^{-6}$ s and $\Delta x=\Delta y=1.4\times 10^{-4}$ m.
@@
In perspective, further studies are planned in ST40 with the two-fold goal of improving the IR thermographic inversion experimentally and numerically. The current IR camera saturation temperature has already been extended from $118$ to $200$, and a detailed cross-comparison with Langmuir probe measurements is underway. The installation of a second IR camera is being considered, for the exploitation of a dual-band adaptor [70] would allow ameliorating any dependence on the uncertain surface emissivity. The commissioning of the dedicated heat-flux testing facility in Tokamak Energy is ongoing, and will pave the way for a validation of FAHF in a controlled environment. Improvements to the numerics of FAHF to handle fast dynamics and impurity surface layers [18, 58] are also taken into account, alongside the implementation of temperature-dependent material properties, and multi-layered PFCs.